OpenStudy (anonymous):

using L'Hopitals lim->0 (cotx-(1/x)

OpenStudy (anonymous):

start with \[\frac{\cos(x)}{\sin(x)}-\frac{1}{x}\] \[=\frac{x\cos(x)-\sin(x)}{x\sin(x)}\] then probably l'hopital